OpenStudy (anonymous):

Square root of (x+deltaX)+ 2 - square root of (x-2)/ delta x

OpenStudy (anonymous):

\[\sqrt{x+\Delta x}+2-\sqrt{\frac{x-2}{\Delta x}}\]??

OpenStudy (anonymous):

no the (delta)x is over both sqaure roots only

OpenStudy (anonymous):

maybe \[\frac{\sqrt{x+\Delta x}+2-\sqrt{x-2}}{\Delta x}\]

OpenStudy (anonymous):

yes but the +2 is also under the square root

OpenStudy (anonymous):

\[\sqrt{(x+\Delta x)+2} - \sqrt{x+2} / \Delta x\]